OpenStudy (anonymous):

consider the population model dp/dt=0.3(1-p/200)(p/50-1)p where p(t) is the population at time t For what values of p is the population in equilibrium? can someone help me with this question...=.="

OpenStudy (anonymous):

derivative is zero ? 0 = 0.3(1-p/200)(p/50-1)p

OpenStudy (anonymous):

p=0,p=50,p=200

OpenStudy (anonymous):

when dp/dt=0 it will be a max or a min.

OpenStudy (anonymous):

i think that i should write the equation more properly since it may give different meaning~~ the exact equation is dp/dt=[0.3][1-(p/200)][(p/50)-1][p] sorry for the inconvenience~~

OpenStudy (anonymous):

i think my answer holds ..

OpenStudy (anonymous):

-.3p^3/1000+1.5p^2/200-.3p=dp/dt

OpenStudy (anonymous):

when the derivative is zero - the population isnt changing.. 0 = 0.3(1-p/200)(p/50-1)p p=0,p=50,p=200

OpenStudy (anonymous):

why ? ana_anne wrote : dp/dt=[0.3][1-(p/200)][(p/50)-1][p]

OpenStudy (anonymous):

-.3p^4/4000+p^3/400-.15p^2+c=p(t)

OpenStudy (anonymous):

nevermind read it wrong. 1-1 duh.

OpenStudy (anonymous):

so they dont need the time, just the equilibrium. 0, 50, 200 are the points of inflection. 0 and 200 are not maxima or minima though. 50 is the equilibrium

OpenStudy (anonymous):

dear coolsector~~ how can i help u~~ i was just trying to make the equation easier to be understand~~=)

OpenStudy (anonymous):

i gave you the solution already..

OpenStudy (anonymous):

the equilibrium occurs at p=50.

OpenStudy (anonymous):

thankz tutordiffeq n coolsector~~ =)

OpenStudy (anonymous):

what about 0,200 ?

OpenStudy (anonymous):

they are points of inflection, not maxima or minima. graph it if you wanna see. i put the integral up

OpenStudy (anonymous):

i dont really understand what's wrong with it .. the derivative is in fact zero there so no change occures

OpenStudy (anonymous):

dear tutordiffeq, what do u mean by u integral it up~~ i cant understand it~~~ =.="